OpenStudy (anonymous):

what is the derivative of 1+ln xy = e^(x-y)

OpenStudy (abb0t):

Are you familiar with the chain rule? Because you want to use that here.

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

Chain rule states: Suppose that you have two functions f(x) and g(x) and they are both differentiable, thus: \[F(x) = (f og)(x)\] and therefore,: \[F'(x)= f'(g(x)) \times g'(x)\] OR similarly: \[\frac{ dy }{ dx } = \frac{ dy }{ du } \frac{ du }{ dx }\]

OpenStudy (abb0t):

It might also help to know the derivative of ln(x): \[\frac{ d }{ dx }\ln(x)=\frac{ 1 }{ x }\] and \[\frac{ d }{ dx }e^x= e^x\]

OpenStudy (abb0t):

In this problem, you are also using product rule. [HINT: xy]

OpenStudy (anonymous):

my main concern is the deriv of e^(x-y). would it be e^(x-y) * y'?

OpenStudy (abb0t):

Yes, that's correct. I'm assuming you are trying to solve for \[\frac{ dy }{ dx }\] So you want to use algebra to rearrange it to get: \[\frac{ dy }{ dx } = f(g(x))\]

OpenStudy (anonymous):

yeah i am

OpenStudy (amoodarya):